# Thermal Modelling of Battery Cells for Optimal Tab and Surface Cooling Control

### v1

@@ I Introduction
Subject to various thermal boundary conditions such as liquid or air convection [4], battery cells can be thermally controlled at different surfaces, including the electrical connection tabs (terminals), cell surfaces, or both [5]. Prior literature studies have investigated the impact of tab and surface cooling methods on battery thermal performance. Hunt et al. and Zhao et al. concluded in [6, 7] that surface cooling can maintain lithium (Li)-ion pouch cells at a lower average temperature under high current rates than tab cooling. In a case study, these authors also demonstrated that using tab cooling rather than surface cooling extended the lifetime of a battery pack by three times. Similar results were achieved in [8, 9] for cylindrical Li-ion cells, where tab cooling was shown to reduce the internal temperature inhomogeneities by about 25%.
@@
It is clear that both surface and tab cooling have their individual strengths and weaknesses. However, to the best of our knowledge, there is no battery control framework in the state-of-the-art literature that systematically investigates the optimal integration of these two cooling methods for advanced battery temperature management.
@@
Thermal equivalent circuit (TEC) models with lumped parameters have been extensively utilized for control-oriented modelling of batteries due to their ease of implementation and computational efficiency [10, 11]. However, lumped parameter TECs can only predict average and surface temperatures and their applicability is limited to cells with small Biot numbers. Physics-based models [12] result in partial differential equations (PDEs) governing the underlying heat diffusion. They can predict the spatially distributed temperature field throughout the cell but are typically implemented via computationally expensive numerical methods, such as finite element (FE) and finite difference (FD) methods, rendering them impractical for control purposes. Spectral methods [13, 14] are alternative numerical methods for finding solutions to PDEs. They belong to the class of weighted residual methods, and unlike FE and FD, they make use of global (rather than local) approximating functions in the discretization of the spatial domain, rendering them computationally efficient. With this benefit in mind, the spectral method based on the Galerkin approach was adopted to develop low-order 2-dimensional (2D) thermal models for cylindrical cells in [15] and for pouch cells in [16]. However, these models do not allow independent and targeted cooling control of battery tabs and surfaces, preventing their use for optimally combining tab and surface cooling for battery thermal management.
@@ II Overview of the PDE-based thermal model
Battery cells of cylindrical and pouch types are widely used for electrical energy storage, as schematized in Figs. 1a-b. Considering the symmetric structure along the radial ($r$) direction of a cylindrical cell, and by assuming the same thermal conductivity in $y$ and $z$ directions of a pouch cell, the temperature distribution of both cell types can be modelled in 2D. For completeness, this section briefly reviews the first principle, PDE-based models for battery thermal dynamics, including mathematical formulations as well as their underlying assumptions and considerations.
@@ II-A Cylindrical cell modelling
For the cylindrical cell, the governing equation for temperature $T(r,z,t)$ at time $t$ and in the position $(r,z)$ can be represented by the following 2D boundary value problem [17]
$\begin{split}&\rho C_{p}\frac{\partial T(r,z,t)}{\partial t}-k_{r}\frac{\partial^{2}T(r,z,t)}{\partial r^{2}}\\ -k_{r}&\frac{\partial T(r,z,t)}{r\partial r}-k_{z}\frac{\partial^{2}T(r,z,t)}{\partial z^{2}}=q(t)\end{split}$ (1)
@@
where $\rho$, $C_{p}$, $k$, $h$, and $T_{\infty}$ represent the volume-average density, volume-average specific heat capacity, thermal conductivity, convection coefficient, and fluid free-stream temperature, respectively. $R_{\text{in}}$ and $R_{\text{out}}$ are the cylindrical cell’s inner and outer radius, and $L$ is the height. The subscripts $s$, $c$, $t$, and $b$ denote the cell’s surface, core, top, and bottom, respectively. Equation (2) is illustrated in Fig. 1b where the convective heat transfer is presumed to occur on the external surfaces, and the heat transfer coefficient and fluid free-stream temperatures may differ for each surface. The cooling occurring in the core area is considered as zero, leading to the corresponding convection coefficient $h_{c}=0$ in (2b).
The cylindrical cell’s multi-layer composition is considered a homogeneous solid with anisotropic thermal conductivities along $r$ and $z$ directions. Due to the spirally wound structure of the electrodes of the cylindrical cell, the thermal conductivity $k_{r}$ would differ from the outer to the inner radius but is assumed constant in this study for simplicity. The heat generation rate $q(t)$ in (1) is assumed to distribute uniformly in space.
For cylindrical cells, the tabs placed on the top and bottom sides can have different sizes. Existing designs make the tab’s area much smaller than the cell’s top/bottom area. It has been demonstrated that these designs of cylindrical cells are not efficient for battery cooling merely from the tabs [8, 18]. In this regard, this work considers the entire area of the top/bottom as the tab, forming the so-called all-tab or tab-less cell design, as illustrated in Figs. 1a and b. Correspondingly, the thermal convection coefficients $h_{t}$ and $h_{b}$ in (2) are affected by the entire top and bottom areas, respectively, in addition to the applied cooling media.
@@ II-B Pouch cell modelling
The thermal model for pouch cells illustrated in Fig. 1c can be derived in a similar fashion as for the cylindrical cell. The thermal conductivity in the $y$ direction is assumed to be identical to that of the $z$ direction because of the stacked structure of the pouch cell. By considering the pouch cell’s multi-layer structure as a homogeneous solid with anisotropic thermal conductivities, $k_{x}$ in the $x$ direction and $k_{y}$ in the $y$ direction, the governing heat equation in Cartesian coordinates is given by the 2D boundary value problem [17]
@@
III Control-oriented battery thermal modeling
@@ III-A Model non-dimensionalization
where $\gamma=\alpha^{2}k_{r}/(1+\tilde{r}+\alpha R_{\text{in}})$. The cooling power applied per unit area is denoted by $u$, and their values at the surface, core, top and bottom sides of the cell are given respectively by
@@
The above cooling power can be considered as the dynamic thermal system’s input. Each of these inputs can be influenced by the coolant’s temperature and flow rate.
@@ III-C Model reduction for the homogeneous solution
where $c_{mn}(t)$ are unknown expansion (solution) coefficients. $\phi_{m}^{\tilde{r}}(\tilde{r})$ is the $m$-th Chebyshev basis function along the dimension $\tilde{r}$, satisfying the boundary conditions (11a)–(11b). $M$ is the number of Chebyshev basis functions, and $m\in\{1,\cdots,M\}$. Along the dimension $\tilde{z}$, $\phi_{n}^{\tilde{z}}(\tilde{z})$ is the $n$-th Chebyshev basis function satisfying the boundary conditions (11c)–(11d), where $n\in\{1,\cdots,N\}$. $M$ and $N$ are tuning parameters depending on the required resolution.
@@
Substituting (17) with the actual form of $\hat{T}_{h}$, we have
@@
Finally, rewriting (18) in matrix-vector notations, it gives
$\begin{split}&\Bigg{\langle}r\Bigg{[}\rho C_{p}\Big{[}\phi_{1}^{\tilde{r}}\phi_{1}^{\tilde{z}},\cdots,\phi_{M}^{\tilde{r}}\phi_{N}^{\tilde{z}}\Big{]}\begin{bmatrix}\frac{d}{dt}c_{11}(t)\\ \vdots\\ \frac{d}{dt}c_{MN}(t)\end{bmatrix}\\ &-\alpha^{2}k_{r}\Bigg{[}\frac{\partial^{2}\phi_{1}^{\tilde{r}}\phi_{1}^{\tilde{z}}}{\partial\tilde{r}^{2}},\cdots,\frac{\partial^{2}\phi_{M}^{\tilde{r}}\phi_{N}^{\tilde{z}}}{\partial\tilde{r}^{2}}\Bigg{]}\begin{bmatrix}c_{11}(t)\\ \vdots\\ c_{MN}(t)\end{bmatrix}\\ &-\gamma\Bigg{[}\frac{\partial\phi_{1}^{\tilde{r}}\phi_{1}^{\tilde{z}}}{\partial\tilde{r}},\cdots,\frac{\partial\phi_{M}^{\tilde{r}}\phi_{N}^{\tilde{z}}}{\partial\tilde{r}}\Bigg{]}\begin{bmatrix}c_{11}(t)\\ \vdots\\ c_{MN}(t)\end{bmatrix}\\ &-\beta^{2}k_{z}\Bigg{[}\frac{\partial^{2}\phi_{1}^{\tilde{r}}\phi_{1}^{\tilde{z}}}{\partial\tilde{z}^{2}},\cdots,\frac{\partial^{2}\phi_{M}^{\tilde{r}}\phi_{N}^{\tilde{z}}}{\partial\tilde{z}^{2}}\Bigg{]}\begin{bmatrix}c_{11}(t)\\ \vdots\\ c_{MN}(t)\end{bmatrix}-q^{*}\Bigg{]},\eta\Bigg{\rangle}=0.\end{split}$ (19)
@@ III-D1 Approximation of $T_{p}$ from a polynomial vector space
where $\otimes$ and $\oplus$ are the outer product and direct sum, respectively. $\tilde{z}$, $\tilde{z}^{2}$, $\tilde{r}$, and $\tilde{r}^{2}$ are the basis vectors of $\mathcal{V}_{rz}$. (20) explicitly couples the $\tilde{r}$ and $\tilde{z}$ dimensions by combining basis functions from one dimension with polynomials in the other. This coupling is essential because the heat generation and boundary conditions can create interactions between $\tilde{r}$ and $\tilde{z}$ that a simpler product space might not capture effectively.
@@
Any function, $f(y)\in\mathcal{V}_{y}$ can be represented as a linear combination of these basis vectors. Specifically, we can write
$f(y)=p_{1}y+p_{2}y^{2},$ (22)
where $p_{1}$ and $p_{2}$ are coefficients to be determined. In many cases, the function $f(y)$ can be complex, and a simple linear combination like (22) may not suffice to represent $f(y)$ accurately. To better approximate $f(y)$, as commonly done in numerical methods like FE analysis and spectral methods, we use a finite-sum approximation with the given basis vectors $\{y,y^{2}\}$. This can be expressed as
$f(y)\approx\sum^{N_{f}}_{i=1}(p_{1,i}y+p_{2,i}y^{2}),$ (23)
where $N_{f}$ is the number of terms in the finite-sum approximation. $p_{1,i}$ and $p_{2,i}$ are the unknown coefficients for each term in the sum, which can be determined using various methods, such as the least squares or spectral methods.
Following the idea of finite-sum approximation introduced in (21)-(23), approximating $T_{p}$ in the vector space $\mathcal{V}_{rz}$ leads to the expansion
$\begin{split}\hat{T}_{p}(\tilde{r},\tilde{z},t)&=\sum^{N}_{n=1}\Big{(}D_{1,n}(\tilde{z},t)\tilde{r}+D_{2,n}(\tilde{z},t)\tilde{r}^{2}\Big{)}\phi_{n}^{\tilde{z}}(\tilde{z})\\ &+\sum^{M}_{m=1}\Big{(}D_{3,m}(\tilde{r},t)\tilde{z}+D_{4,m}(\tilde{r},t)\tilde{z}^{2}\Big{)}\phi_{m}^{\tilde{r}}(\tilde{r}),\end{split}$ (24)
@@ III-D2 Determination of the expansion coefficients
$R_{s}=h_{s}\hat{T}_{p}-\alpha k_{r}\frac{\partial\hat{T}_{p}}{\partial\tilde{r}}-u_{s}\neq 0,\>\>at\>\>\tilde{r}=1.$ (25)
@@
$\langle R_{s},\phi_{i}^{\tilde{z}}\rangle_{\tilde{z}}=\Bigg{\langle}\bigg{[}h_{s}\hat{T}_{p}-\alpha k_{r}\frac{\partial\hat{T}_{p}}{\partial\tilde{r}}-u_{s}\bigg{]},\phi_{i}^{\tilde{z}}\Bigg{\rangle}_{\tilde{z}}=0,\>\>at\>\>\tilde{r}=1,$ (26)
@@
Then, we adopt a similar process for the other vertical side (i.e., the core) and the horizontal sides (i.e., the top and bottom sides) of the battery cell with the boundary conditions formulated in (13b)–(13d). Consequently, we have
@@
$\sum^{M}_{n=1}$ $\Big{[}D_{3,m}(h_{t}-\beta k_{z})+D_{4,m}(h_{t}-2\beta k_{z})\Big{]}\Big{\langle}\phi_{m}^{\tilde{r}},\phi_{j}^{\tilde{r}}\Big{\rangle}_{\tilde{r}}$
@@
$\sum^{M}_{n=1}$ $\Big{[}D_{3,m}(h_{b}-\beta k_{z})+D_{4,m}(-h_{b}+2\beta k_{z})\Big{]}\Big{\langle}\phi_{m}^{\tilde{r}},\phi_{j}^{\tilde{r}}\Big{\rangle}_{\tilde{r}}$
@@
where the matrices for the vertical sides $\Phi_{v}(\tilde{z})\in\mathcal{R}^{N\times N}$ and the horizontal sides $\Phi_{h}(\tilde{r})\in\mathcal{R}^{N\times N}$ are given by $\Phi_{v}=\langle\phi_{n}^{\tilde{z}},\phi_{i}^{\tilde{z}}\rangle_{\tilde{z}}$ and $\Phi_{h}=\langle\phi_{m}^{\tilde{r}},\phi_{j}^{\tilde{r}}\rangle_{\tilde{r}}$, respectively. The source terms for the vertical sides $S_{v}(\tilde{z})\in\mathcal{R}^{N}$ and the horizontal sides $S_{h}(\tilde{r})\in\mathcal{R}^{M}$ are given by $S_{v}=\langle 1,\phi_{i}^{\tilde{z}}\rangle_{\tilde{z}}$ and $S_{h}=\langle 1,\phi_{j}^{\tilde{r}}\rangle_{\tilde{r}}$, respectively. $D_{1}$, $D_{2}$, $D_{3}$, and $D_{4}$ are vectors of the unknown coefficients, defined explicitly as

### v2

@@ I Introduction
Subject to various thermal boundary conditions such as liquid or air convection [4], battery cells can be thermally controlled at different surfaces, including the electrical connection tabs (terminals), cell surfaces, or both [5]. Prior literature studies have investigated the impact of tab and surface cooling methods on battery thermal performance. Hunt et al. and Zhao et al. concluded in [6, 7] that surface cooling is more effective at maintaining lower average temperatures in lithium (Li)-ion pouch cells, especially under high current rates, due to the larger cooling area. In a case study, these authors also demonstrated that tab cooling extended the lifetime of a battery pack by a factor of three, primarily due to the more uniform temperature distribution it provides. In a follow-up study, Dondelewski et al. compared both methods in [8] and found that surface cooling resulted in lower average temperatures and slower degradation but reduced usable capacity. Conversely, tab cooling produced higher average temperatures and faster degradation, yet allowed for greater usable capacity due to more effective thermal distribution. Similar results were achieved for cylindrical Li-ion cells in [9, 10], where tab cooling was shown to reduce internal temperature inhomogeneities by approximately 25%, though it resulted in a higher average cell temperature. Furthermore, [11] reported that surface cooling causes a faster drop in temperature, driven by the larger lateral surface area available for heat transfer. However, tab cooling offered improved thermal uniformity, with a recorded temperature difference of 1.9°C, compared to 3.8°C under surface cooling.
@@
It is clear that both surface and tab cooling methods offer distinct advantages; surface cooling can enable faster heat removal and lower average cell temperatures, while tab cooling enhances temperature uniformity and thermal balance. However, to the best of our knowledge, there is currently no battery control framework in the state-of-the-art literature that systematically investigates the optimal integration of these two cooling methods for advanced battery thermal management.
@@
Thermal equivalent circuit (TEC) models with lumped parameters have been extensively utilized for control-oriented modelling of batteries due to their ease of implementation and computational efficiency [12, 13]. However, lumped parameter TECs can only predict average and surface temperatures and their applicability is limited to cells with small Biot numbers. Physics-based models [14] result in partial differential equations (PDEs) governing the underlying heat diffusion. They can predict the spatially distributed temperature field throughout the cell but are typically implemented via computationally expensive numerical methods, such as finite element (FE) and finite difference (FD) methods, rendering them impractical for real-time control purposes. Spectral methods [15, 16] are alternative numerical methods for finding solutions to PDEs. They belong to the class of weighted residual methods, and unlike FE and FD, they make use of global (rather than local) approximating functions in the discretization of the spatial domain, rendering them computationally efficient. With this benefit in mind, the spectral method based on the Galerkin approach was adopted to develop low-order 2-dimensional (2D) thermal models for cylindrical cells in [17] and for pouch cells in [18]. However, these models do not allow independent and targeted cooling control of battery tabs and surfaces, preventing their use for optimally combining tab and surface cooling for battery thermal management.
@@ II Overview of the PDE-based thermal model
Battery cells of cylindrical and pouch types are widely used for electrical energy storage, as schematized in Figs. 1(a)-(c). Considering the symmetric structure along the radial ($r$) direction of a cylindrical cell, and by assuming the same thermal conductivity in $y$- and $z$-directions of a pouch cell, the temperature distribution of both cell types can be modelled in 2D. For completeness, this section briefly reviews the first principle PDE-based models for battery thermal dynamics, including mathematical formulations as well as their underlying assumptions and considerations.
@@ II-A Cylindrical cell modelling
For the cylindrical cell, the governing equation for temperature $T(r,z,t)$ at time $t$ and position $(r,z)$ can be represented by the following 2D boundary value problem [19]:
$\begin{split}&\rho c_{p}\frac{\partial T(r,z,t)}{\partial t}-k_{r}\frac{\partial^{2}T(r,z,t)}{\partial r^{2}}\\ -k_{r}&\frac{\partial T(r,z,t)}{r\partial r}-k_{z}\frac{\partial^{2}T(r,z,t)}{\partial z^{2}}=q(t),\end{split}$ (1)
@@
where $\rho$, $c_{p}$, $k$, $h$, and $T_{\infty}$ represent the volume-average density, volume-average specific heat capacity, thermal conductivity, convection coefficient, and fluid free-stream temperature, respectively. $R_{\text{in}}$ and $R_{\text{out}}$ are the cylindrical cell’s inner and outer radius, and $L$ is the length. The subscripts $s$, $c$, $t$, and $b$ denote the cell’s surface, core, top, and bottom, respectively. Equation (2) is illustrated in Fig. 1(b), where the convective heat transfer is presumed to occur on the external surfaces, and the heat transfer coefficient and fluid free-stream temperatures may differ for each surface. The cooling occurring in the core area is considered as zero, leading to the corresponding convection coefficient $h_{c}=0$ in (2b).
The cylindrical cell’s multi-layer composition is considered a homogeneous solid with anisotropic thermal conductivities along $r$- and $z$-directions. Due to the spirally wound structure of the electrodes of the cylindrical cell, the thermal conductivity $k_{r}$ would differ slightly from the inner to the outer radius but is assumed constant in this study for simplicity. The heat generation rate $q(t)$ in (1) is assumed to be uniformly distributed in space.
For cylindrical cells, the tabs placed on the top and bottom sides can have different sizes. Existing designs make the tab’s area much smaller than the cell’s top/bottom area. It has been demonstrated that these designs of cylindrical cells are not efficient for battery cooling merely from the tabs [9, 20, 21]. In this regard, this work considers the entire area of the top/bottom as the tab, forming the so-called all-tab or tab-less cell design, as illustrated in Figs. 1a and b. Correspondingly, the thermal convection coefficients $h_{t}$ and $h_{b}$ in (2) are affected by the entire top and bottom areas, respectively, in addition to the applied cooling media.
@@ II-B Pouch cell modelling
The thermal model for pouch cells illustrated in Fig. 1(c) can be derived in a similar fashion as for the cylindrical cell. The thermal conductivity in the $y$-direction is assumed to be identical to that of the $z$-direction because of the stacked structure of the pouch cell. By considering the pouch cell’s multi-layer structure as a homogeneous solid with anisotropic thermal conductivities $k_{x}$ in the $x$-direction and $k_{y}$ in the $y$-direction, the governing heat equation in Cartesian coordinates is given by the 2D boundary value problem [19]
@@
III Control-oriented battery thermal modelling
@@ III-A Model non-dimensionalization
where $\gamma=\alpha^{2}k_{r}/(1+\tilde{r}+\alpha R_{\text{in}})$. The cooling power applied per unit area is denoted by $u$, and their values at the surface, core, top and bottom sides of the cell are given, respectively, by
@@
The above cooling power can be considered as the dynamic thermal system’s input. The coolant’s temperature and flow rate can influence each of these inputs. Furthermore, (8) can be integrated with a flow model to better capture the convective cooling effects and dynamic interactions between the coolant and the battery system under various cooling scenarios.
@@ III-C Model reduction for the homogeneous solution
where $c_{mn}(t)$ are unknown expansion (solution) coefficients. $\phi_{m}^{\tilde{r}}(\tilde{r})$ is the $m$-th Chebyshev basis function along the dimension $\tilde{r}$, satisfying the boundary conditions (11a)–(11b). $M$ is the number of Chebyshev basis functions, and $m\in\{1,\cdots,M\}$. Along the dimension $\tilde{z}$, $\phi_{n}^{\tilde{z}}(\tilde{z})$ is the $n$-th Chebyshev basis function satisfying the boundary conditions (11c)–(11d), where $n\in\{1,\cdots,N\}$. $M$ and $N$ are tuning parameters depending on the required accuracy.
@@
Substituting $\hat{T}_{h}$ in (17) with its actual form in (14), we have
@@
Finally, rewriting (18) in matrix-vector notations, this gives
$\Bigg{\langle}r\Bigg{[}\rho c_{p}\Big{[}\phi_{1}^{\tilde{r}}\phi_{1}^{\tilde{z}},\cdots,\phi_{M}^{\tilde{r}}\phi_{N}^{\tilde{z}}\Big{]}\begin{bmatrix}\frac{d}{dt}c_{11}(t)\\ \vdots\\ \frac{d}{dt}c_{MN}(t)\end{bmatrix}$
$-\alpha^{2}k_{r}\Bigg{[}\frac{\partial^{2}\phi_{1}^{\tilde{r}}\phi_{1}^{\tilde{z}}}{\partial\tilde{r}^{2}},\cdots,\frac{\partial^{2}\phi_{M}^{\tilde{r}}\phi_{N}^{\tilde{z}}}{\partial\tilde{r}^{2}}\Bigg{]}\begin{bmatrix}c_{11}(t)\\ \vdots\\ c_{MN}(t)\end{bmatrix}$
$-\gamma\Bigg{[}\frac{\partial\phi_{1}^{\tilde{r}}\phi_{1}^{\tilde{z}}}{\partial\tilde{r}},\cdots,\frac{\partial\phi_{M}^{\tilde{r}}\phi_{N}^{\tilde{z}}}{\partial\tilde{r}}\Bigg{]}\begin{bmatrix}c_{11}(t)\\ \vdots\\ c_{MN}(t)\end{bmatrix}$
$-\beta^{2}k_{z}\Bigg{[}\frac{\partial^{2}\phi_{1}^{\tilde{r}}\phi_{1}^{\tilde{z}}}{\partial\tilde{z}^{2}},\cdots,\frac{\partial^{2}\phi_{M}^{\tilde{r}}\phi_{N}^{\tilde{z}}}{\partial\tilde{z}^{2}}\Bigg{]}\begin{bmatrix}c_{11}(t)\\ \vdots\\ c_{MN}(t)\end{bmatrix}-q^{*}\Bigg{]},\eta\Bigg{\rangle}=0.$ (19)
@@ III-D1 Approximation of $T_{p}$ from a polynomial vector space
where $\otimes$ and $\oplus$ are the outer product and direct sum, respectively. $\tilde{z}$, $\tilde{z}^{2}$, $\tilde{r}$, and $\tilde{r}^{2}$ are the basis vectors of $\mathcal{V}_{rz}$. (20) explicitly couples the $\tilde{r}$- and $\tilde{z}$-dimensions by combining basis functions from one dimension with polynomials in the other. This coupling is essential because the heat generation and boundary conditions can create interactions between $\tilde{r}$ and $\tilde{z}$ that a simpler product space might not capture effectively.
@@
Any function, $f(y)\in\mathcal{V}_{y}$ can be represented as a linear combination of these basis vectors. Specifically, we can write any such $f(y)$ using a finite-sum approximation
$f(y)\approx\sum^{2}_{i=1}p_{i}y^{i},$ (22)
where $p_{i}$ are the unknown coefficients, which can be determined using various methods, such as the least squares or spectral methods.
Following the idea of finite-sum approximation introduced in (21)-(22), approximating $T_{p}$ in the vector space $\mathcal{V}_{rz}$, i.e., in two space dimensions, leads to the expansion
$\hat{T}_{p}(\tilde{r},\tilde{z},t)$ $=\sum^{N}_{n=1}\Big{(}D_{1,n}(\tilde{z},t)\tilde{r}+D_{2,n}(\tilde{z},t)\tilde{r}^{2}\Big{)}\phi_{n}^{\tilde{z}}(\tilde{z})$
$+\sum^{M}_{m=1}\Big{(}D_{3,m}(\tilde{r},t)\tilde{z}+D_{4,m}(\tilde{r},t)\tilde{z}^{2}\Big{)}\phi_{m}^{\tilde{r}}(\tilde{r}),$ (23)
@@ III-D2 Determination of the expansion coefficients
$R_{s}=h_{s}\hat{T}_{p}-\alpha k_{r}\frac{\partial\hat{T}_{p}}{\partial\tilde{r}}-u_{s}\neq 0,\>\>\mbox{at}\>\>\tilde{r}=1.$ (24)
@@
$\langle R_{s},\phi_{i}^{\tilde{z}}\rangle_{\tilde{z}}=\Bigg{\langle}\bigg{[}h_{s}\hat{T}_{p}-\alpha k_{r}\frac{\partial\hat{T}_{p}}{\partial\tilde{r}}-u_{s}\bigg{]},\phi_{i}^{\tilde{z}}\Bigg{\rangle}_{\tilde{z}}=0,\>\>\text{at}\>\>\tilde{r}=1,$ (25)
@@
Then, we adopt a similar process for the other vertical side (i.e., the core) and the horizontal sides (i.e., the top and bottom sides) of the battery cell with the boundary conditions formulated in (13b)–(13d). Consequently, we get
@@
$\sum^{M}_{m=1}$ $\Big{[}D_{3,m}(h_{t}-\beta k_{z})+D_{4,m}(h_{t}-2\beta k_{z})\Big{]}\Big{\langle}\phi_{m}^{\tilde{r}},\phi_{j}^{\tilde{r}}\Big{\rangle}_{\tilde{r}}$
@@
$\sum^{M}_{m=1}$ $\Big{[}D_{3,m}(h_{b}-\beta k_{z})+D_{4,m}(-h_{b}+2\beta k_{z})\Big{]}\Big{\langle}\phi_{m}^{\tilde{r}},\phi_{j}^{\tilde{r}}\Big{\rangle}_{\tilde{r}}$
@@
where the matrices for the vertical sides $\Phi_{v}(\tilde{z})\in\mathcal{R}^{N\times N}$ and the horizontal sides $\Phi_{h}(\tilde{r})\in\mathcal{R}^{N\times N}$ are given by $\Phi_{v}=\langle\phi_{n}^{\tilde{z}},\phi_{i}^{\tilde{z}}\rangle_{\tilde{z}}$ and $\Phi_{h}=\langle\phi_{m}^{\tilde{r}},\phi_{j}^{\tilde{r}}\rangle_{\tilde{r}}$, respectively. The source terms for the vertical sides $S_{v}(\tilde{z})\in\mathcal{R}^{N}$ and the horizontal sides $S_{h}(\tilde{r})\in\mathcal{R}^{M}$ are given by $S_{v}=\langle 1,\phi_{i}^{\tilde{z}}\rangle_{\tilde{z}}$ and $S_{h}=\langle 1,\phi_{j}^{\tilde{r}}\rangle_{\tilde{r}}$, respectively. $D_{1}$, $D_{2}$, $D_{3}$, and $D_{4}$ are vectors of the unknown coefficients, explicitly defined as